Perform the following modular arithmetic by applying the addition/subtraction rule.

(a) (9+15) mod 7

(b) (15-23) mod 5

(c) (-47 x (5+1)) mod 5


1
Expert's answer
2021-11-18T11:49:32-0500

Let us perform the following modular arithmetic by applying the addition/subtraction rule.


(a) "(9+15) \mod 7=(9\mod 7)+(15 \mod 7)=(2\mod 7)+(1 \mod 7)=3 \mod 7"


(b) "(15-23) \mod 5=(15 \mod 5)-(23 \mod 5)=(0 \mod 5)-(3 \mod 5)\\"

"=-3 \mod 5=2 \mod 5"


(c) "(-47 \cdot (5+1)) \mod 5
=(-47 \mod 5)\cdot ((5+1) \mod 5)"

"=(3 \mod 5)\cdot (1 \mod 5)
=3 \mod 5"